if o is the origin and a(1,2) ,Bb(2,3) then circum radius of triangle OAB in sq. units

Grade:11

1 Answers

Vikas TU
5 years ago
Dear Student,
OA=b=sqrt((1-0)^2 +(2-0)^2) =sqrt(5)
similarly OB=a=sqrt(13) and AB=o=sqrt(2)
using Heron’s formula, area =sqrt(s(s-a)(s-b)(s-c)):
 area of triangle OAB=A=1/2sq unit
Circumradius R=oab/4A= sqrt(65/2) units [Ans]
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)